Problem A.1


The graph below is made of three line segments:


-1 1 2 3 4 5 6 7 8 9 10 11 12


1


2


3


4


y


x


f(x)


g(x)


h(x)


The segments correspond to the following three functions:


f(x) = x − 2, g(x) = p


4 − (x − 6)2 + 2, h(x) = x − 6


Find the total length L of the graph between x = 2 and x = 10.

1
Expert's answer
2021-11-22T17:17:11-0500

(i) Length of f(x)

Given,

"f(x)=x-2, \quad for 2 \leq x \leq 4"

The length would be the length of hypotenuse of right angled isosceles triangle of side length =2 units

 "\begin{aligned}

&\Rightarrow h^{2}=b^{2}+h^{2} \\

&\Rightarrow h^{2}=2^{2}+2^{2} \\

&\Rightarrow h^{2}=8 \\

&\Rightarrow h=2 \sqrt{2}

\end{aligned}"

Length of graph contributed by "f(x)=2 \sqrt{2}"

(ii) Length of g(x)

Given,

"\begin{aligned}

&g(x)=\sqrt{4-(x-6)^{2}}+2, \quad \text { for } 4 \leq x \leq 8 \\

&\Rightarrow g(x)-2=\sqrt{4-(x-6)^{2}} \\

&\Rightarrow(g(x)-2)^{2}=4-(x-6)^{2}\\
&\Rightarrow (x-6)^{2}+(g(x)-2)^{2}=2^2
\end{aligned}"

So, g(x) is a circle with centre at (6,2) and radius = 2 units

Length contributed by g(x) would be circumference of semicircle

Length contributed by g(x) = "\pi r=2\pi"


(iii) Length of "\mathrm{h}(\mathrm{x})"

Given,

"f(x)=x-6, \quad for \ 8 \leq x \leq 10"

The length would be the length of hypotenuse of right angled triangle of base = height =2 units

"\begin{aligned}

&\Rightarrow h^{2}=b^{2}+h^{2} \\

&\Rightarrow h^{2}=2^{2}+2^{2} \\

&\Rightarrow h^{2}=8 \\

&\Rightarrow h=2 \sqrt{2}

\end{aligned}"  

Length of graph contributed by "h(x)=2 \sqrt{2}"

Total length of graph

"\begin{aligned}

&L=\text { length contributed by } f(x)+\text { length contributed by } g(x)+\text { length contributed by } h(x) \\

&\Rightarrow L=2 \sqrt{2}+2 \pi+2 \sqrt{2} \\

&\Rightarrow L=4 \sqrt{2}+2 \pi

\end{aligned}"